Loss Amount ($) Probability (%) 600 1,200 3,000 7,000 (a) What is Alfies expected loss amount? (b) What is Alfies standard
Answer #1

(A) Expected loss is the mean value

Hr= [* P(2) (0 * 0.15) + (600 * 0.35) + (1200*0.28) + (3000 * 0.14) + (7000 * 0.08) = 0 + 210 + 336 + 420 + 560 = 1526

(B) standard deviation calculation

σε = Σ2 και P(x)] - με = [(02 και 0.15) + (6002 και 0.35) + (12002 και 0.28) + (30002 και 0.14) + (70002 και 0.08)] – 15262 =

(C) Coefficient of variation = [(standard deviation)/mean]*100

=(1838.620/1526)*100

= 120.49%